. Please show how to solve A package delivery service uses vans and employees to deliver...

.

Please show how to solve

A package delivery service uses vans and employees to deliver the maximum number of packages given a fixed budget. The last van added 600 packages to total output, while the last employee added 500 packages. If vans cost $400 per week and employees earn $300 per week, the firm a. could deliver more packages with the same budget by using more employees and fewer vans. b. could deliver more packages with the same budget by using more vans and fewer employees. c. should use more vans and fewer employees because the last dollar spent on vans added more to total output than the last dollar spent on employees. d. is delivering the maximum number of packages given the fixed budget. e. both b and c
Answer #1

Marginal product from lash van=600 package

marginal product from last employee=500 package

price of vans=400 and price of employee=300

Since MPv/Pv=600/400=1.5

MPe/Pe=500/300=1.667

Since MPe/Pe>MPv/Pv leading to inefficiency.

Thus hire more employee and less van

Thus ans is A